OpenStudy (anonymous):

in ABC, m B is three times as large as m A. an exterior angle at C measures 140. find m A

OpenStudy (anonymous):

let mA be 'x' then mB=3x now we know exterior angle = sum of two interior opposite angle hence by the problem 140=x+3x 140 = 4x or x=140/4 or x=35 thus <A =35